OpenStudy (anonymous):

multiply and simplify: (7-radical 3) (9-radical 12)

OpenStudy (anonymous):

shadow that isnt right.

OpenStudy (anonymous):

oh i thought he was trying to say * not subtract....

OpenStudy (anonymous):

63 - 6 -9rad3 -7rad12 = 30 - 9rad3 -14rad3 = 30 - 23rad3

OpenStudy (anonymous):

\[(7-\sqrt{3})(9-2\sqrt{3})\] is a start

OpenStudy (anonymous):

sorry its 57 - 23rad3

OpenStudy (anonymous):

then do the 4 necessary multiplications and you will get it.

OpenStudy (anonymous):

\[7\times 9 -7\times 2\sqrt{3}-9\times \sqrt{3}+2\sqrt{3}\times \sqrt{3}\]

myininaya (myininaya):

\[7*9-7*\sqrt{12}-\sqrt{3}*9+\sqrt{3}\sqrt{12}=63-7\sqrt{12}-9\sqrt{3}+\sqrt{36}\] =\[63-7*\sqrt{3*4}-9\sqrt{3}+6=63-6-7\sqrt{4}\sqrt{3}-9\sqrt{3}=57-7*2\sqrt{3}-9\sqrt{3}\] =\[57-14\sqrt{3}-9\sqrt{3}=57-23\sqrt{3}\]

OpenStudy (anonymous):

\[63-14\sqrt{3}-9\sqrt{3}+6\] \[69-23\sqrt{3}\]

myininaya (myininaya):

lol

OpenStudy (anonymous):

ok i get 69 and not 57 and i am sticking with my answer

OpenStudy (radar):

I like that much better.

myininaya (myininaya):

satellite is right

OpenStudy (anonymous):

with all gory details worked out, but if any step is not clear let me know

myininaya (myininaya):

jimmy and i got the same wrong answer lol

OpenStudy (anonymous):

yes - its 63 + 6 not 63-6 - satellite is correct

OpenStudy (anonymous):

\[\color{pink}{\text{very very sleep....}}\]

myininaya (myininaya):

yes i probably need sleep

myininaya (myininaya):

i'm gonna get my hours fix today i will not sleep until 8 pm

OpenStudy (anonymous):

then sleep for 24 hours.

myininaya (myininaya):

i cant

OpenStudy (radar):

Another concensus!

OpenStudy (anonymous):

i got 2 wrong answers - lol!

myininaya (myininaya):

lol

OpenStudy (anonymous):

From Mathematica:\[N\left[\left(7-\sqrt{3}\right)\left(9-\sqrt{12}\right),20\right]=29.162831425915822249 \]\[N\left[69-23 \sqrt{3},20\right]=29.162831425915822249 \]\[N\left[57-23\sqrt{3},20\right]=17.162831425915822249 \]

OpenStudy (anonymous):

well whats 12 between friends

myininaya (myininaya):

yes i already seen my mistake for some reason i did 63-6 when we had 63+6

OpenStudy (anonymous):

yeah well i looked at 63 and made it 36 (dyxlexia???}

myininaya (myininaya):

yes i blame it on that too